A potential difference of 5.60 V will be applied to a 39.00 m length of 18-gauge copper wire (diameter = 0.0400inches). Calculate the current. Tries 0/20 Calculate the current density. Tries 0/20 Calculate the electric field. Tries 0/20 Calculate the rate at which thermal energy will appear in the wire.
Answer #1

d = 0.04 in = 1.016* 10^-3 m

a)

resistance of the wire

R = rho * L/( pi /4 d^2)

R = 1.7* 10^-8* 39 / ( 3.14/4* ( 1.016* 10^-3)^2)

R = 0.8182 ohm

current through the wire

i = V/R = 5.6 / 0.8182 = 6.844 A

========

b)

current density

J = i/ (pi/4 d^2) = 6.844 / ( 3.14 / 4* (1.016* 10^-3)^2)

J = 8.446* 10^6 A/m^2

======

c)

P = i^2 R = 6.844^2* 0.8182 = 38.325 W

======

Comment before rate in case any doubt, will reply for sure.. goodluck